OpenStudy (anonymous):

can someone help me log5(x-2)+log8(x-4)=log6(x-1)

hartnn (hartnn):

is that \(\Large \log_5(x-2)\) or \(\Large \log5(x-2)\) ?

OpenStudy (anonymous):

log5(x-2)

OpenStudy (anonymous):

I know how to resolve the problem but thank you for your help:D

hartnn (hartnn):

oh, the 2nd one. ok. we use, the property that \(\log A+\log B=\log(AB)\) so what about \(\log5(x-2)+\log8(x-4)=...\) oh cool!

hartnn (hartnn):

if you want me to verify then post the final answer, i'll verify :)

OpenStudy (anonymous):

ok